OFFER TO PURCHASE REAL ESTATE

From the Of�ce of:

Leonard J. Colwell Real Estate, Inc. 4 Vinebrook Road Lexington, MA 02173

To: Foo and Bar Baz 33 Quux Lane Arlington, MA 02174

Date: September 4, 1993

The property herein referred to is identi�ed as follows:

33 Quux Lane, Arlington, MA 02174, Book xxx page yyy S.M.R.D.

Special provisions (if any) re �xtures, appliances, etc.: including refrigerator

I hereby offer to buy said property, which has been offered to me by Leonard J. Colwell Real Estate, Inc., Carole Shoupe as my broker, and Fly-By-Night Real Estate under the following terms and conditions:

1. I will pay therefore � xxx,xxx of which

(a) � 750 is paid herewith as a deposit to bind this Offer, (Check, subject to collection)

(b) � 10,000 is to be paid as an additional deposit upon the execution of the Purchase and Sale Agreement provided for below,

(c) � yyy,yyy is to be paid at the time of delivery of the Deed in cash, or by certi�ed, cashier’s treasurer’s or bank check(s),

(d) � xxx,xxx Total Purchase Price

2. This Offer is good until 11:00 P.M. on April 28, 1993 at or before which time a copy hereof shall be signed by you, the Seller and your (husband) (wife), signifying acceptance of this Offer, and returned to me forthwith, otherwise this Offer shall be considered as rejected and the money deposited herewith shall be returned to me forthwith.

3. The parties hereto shall, on or before 5 P.M. on May 20, 1993 execute a Purchase and Sale Agreement satisfactory to Buyer and Buyers’ attorney, which, when executed, shall be the agreement between the parties hereto. If agreement on the content of a Purchase and Sale Agreement cannot be reached this offer shall be null and void, and without recourse to either party, and all deposits shall be returned to Buyer.

4. A good and suf�cient Deed, conveying a good and clear record and marketable title shall be delivered at 12:00 Noon on July 9, 1993 at the appropriate Registry of Deeds, unless some other time and place are mutually agreed upon in writing.

5. If I do not ful�ll my obligations under this Offer, the above mentioned deposit shall forthwith become your property without recourse to either party. Said deposit shall be held by Leonard J. Colwell Real Estate, Inc. as escrow agent subject to the terms hereof provided however that in the event of any disagreement between the parties, the escrow agent may retain said deposit pending instructions mutually given by the parties. A similar provision shall be included in the Purchase and Sale Agreement with respect to any deposits held under its terms.

6. (INSPECTIONS) This offer is subject to Buyer obtaining a satisfactory building, pest, radon, and lead paint inspection report before May 19, 1993. In the event that defects are discovered in the course of the aforementioned inspection(s), the Buyer shall supply to the Seller(s) the relevant portions of the inspector(s) written report. Said defects shall be corrected at the Seller’(s) expense, provided that the Seller(s) shall so agree; otherwise this agreement shall become null and void, without recourse to either party, and all deposits shall be returned to Buyer.

7. (FINANCING CONTINGENCY) This offer is subject to the Buyer’s securing a written commitment from a recognized lending institution for a direct reduction mortgage loan of not more than � lll,ddd. If despite Buyer’s diligent efforts a commitment for such loan cannot be obtained on or before June 25, 1993 the Buyer may terminate this agreement by written notice to the seller and/or the Broker(s), as agent for the Seller(s), prior to the expiration of such time, whereupon any payment made under this agreement shall be forthwith refunded and all other obligations of the parties hereto shall cease and this agreement shall be void without recourse to the parties hereto. In no event will the Buyer be deemed to have used diligent efforts to obtain such commitment unless the Buyer submits to a lending institution a complete mortgage loan application on or before May 24, 1993. The Buyer recognizes that lending institutions may require the payment of an application fee, credit report charge, origination fee, commitment fee (points), and other charges as part of the loan application process. The demand by the lending institution of these charges does not relieve the Buyer of the obligation to apply for said commitment. The Buyer shall also keep the Broker(s) informed as to the banks(s) involved and will so inform the bank(s) as to the name(s) of the Broker(s) involved and so instruct the bank(s) to inform said Broker(s) upon their request as to the progress of the application and loan through closing.

8. (AGENCY DISCLOSURE) In accordance with Regulation 254 of the Code of Massachuesetts Regulations, Section 2.05 (15), the undersigned prospective Purchaser/Buyer, herwith acknowledges that he has been duly informed, prior to proposing to purchase this property, that all Brokers/Salespersons represent the seller, not the buyer, in the marketing, negoiating and sales of property, unless otherwise disclosed. However, the Broker or Salesperson has an ethical and legal obligation to show honesty and fairness to the buyer in all transactions.

9. Time is of the essence hereof.

10. The initialed riders, if any, attached hereto are incorporated herein by reference. Additional terms and conditions, if any:

NOTICE: This is a legal document that creates binding obligations. If not understood, consult an attorney.

WITNESS my hand and seal SIGNED

Buyer  John T. Kohl

Parlmont Park, Apt � 111 N. Billerica, MA 01862

(508) 663-3453

This Offer is hereby accepted upon the foregoing terms and conditions at A.M./P.M. on 1993.

WITNESS my (our) hand(s) and seal(s).

Seller (or spouse) Seller

Received from John Kohl Buyer the sum of �750 as a deposit under the terms and conditions of above Offer, to be held by Leonard J. Colwell Real Estate, Inc. as escrow agent.

Agent for Seller